i have 2 questions first is

1). Convert 3.25x10^5 grams pb into moles.


2). convert 31 mo AI into AI



1
Expert's answer
2021-02-22T05:57:05-0500

A)Mass= 3.25×10"^5"

Molar mass= 207

Mole=mass/molar mass

Mole= 3.25×10^5/207

Mole= 0.015×10"^5"

B) not readable